OpenStudy (anonymous):

What is the height of each ball on the fifth bounce (i.e., Height 6)? Use the geometric sequence formula, an = a1rn – 1 and show your work.

OpenStudy (anonymous):

Ball 1 Description Ball 2 Description Ball 3 Description Height 1 (start) 3 ft 3 ft 3 ft Height 2 2.4 1.9 0.9 Height 3 1.9 1.2 0.3 Height 4 1.5 0.8 0.1 Height 5 1.2 0.5 0.02

OpenStudy (anonymous):

visit this site and scroll down.. http://inquus.org/updates/53a47d88e4b0a819ab10438f

OpenStudy (anonymous):

thank you but that didn't help me

OpenStudy (anonymous):

1/ find r by pick any of them. I pick \(a_5\) to calculate r . Apply formula \(a_5=a_1r^4\), we have \(r=\sqrt[4]{\dfrac{a_5}{a_1}}=\sqrt[4]{\dfrac{1.2}{3}}=0.795\) 2/ Now, we have r, just apply the formula to find \(a_6\) for each of balls.